KTS is trying to estimate the first-year operating cash flow. The financial staff has given you the following information:

Incremental revenue = $8

Depreciation = $2

Interest Expense = $1

Marginal tax rate = 50%

What is the operating cash flow in year 1?
Answer #1

OCF = (revenues-depr-interest expense)*(1-tax rate)+depr = (8-2-1)*(1-0.5)+2 = 4.5

phpfU8RLb.png